OpenStudy (ariannamoniquee97):

PLEASE HELP! :)) An artist makes three types of ceramic statues (large, medium, and small) at a monthly cost of $890 for 190 statues. The manufacturing costs for the three types are $6, $5, and $4. If the statues sell for $24, $16, and $13, respectively, how many of each type should be made to produce $3,060 in monthly revenue?

OpenStudy (robtobey2):

Let L, M and S be the number of large, medium and small statues respectively. From the problem statement: { L + M + S = 190, 6 L + 5 M + 4 S = 890, 24 L + 16 M + 13 S = 3060 } Solve the above for L, M and S . {L = 40, M = 50, S = 100}